a car is traveling 3m/s decelerates at 7m/s*2. How long is it before it comes to the rest?


1
Expert's answer
2020-09-22T15:35:20-0400

The velocity of the car which moves with the constant (and negative, because the car is decelerating) acceleration is:


"v = v_0 - at"

where "v_0 = 3m\/s" is the initial speed, "t" is the time and "a = 7m\/s^2" is the acceleration.

At rest "v = 0". Thus:


"0 = v_0 - at\\
t = \dfrac{v_0}{a} = \dfrac{3}{7}s\approx 0.43s"

Answer. 3/7s or 0.43s.